OpenStudy (anonymous):

f(x)=-2x+4 and g(x)=-6x-7. Find f(x)-g(x)

OpenStudy (dan815):

-2x+4 - (-6x-7)=4x+11

OpenStudy (mimi_x3):

f(x)-g(x)=(-2x+4)-(-6x-7) =-2x+4+6x+7 =4x+11

OpenStudy (dan815):

phew i beat her just in time

OpenStudy (mimi_x3):

fu

Nnesha (nnesha):

coc please

OpenStudy (mimi_x3):

wingspan please?????? Watch ur language nnesha

Nnesha (nnesha):

my language O_O what i said ??? is totally positive

Nnesha (nnesha):

its**